# Geometry of efficient weight vectors

### v1

@@ Abstract
Pairwise comparison matrices and the weight vectors obtained from them are important concepts in multi-criteria decision making. A weight vector corresponding to a pairwise comparison matrix is called Pareto efficient if the approximation of the matrix elements by the weight ratios cannot be improved for any element of the matrix without worsening it for an other element. The aim of this paper is to show the geometrical properties of the set of Pareto efficient weight vectors for $4\times 4$ pairwise comparison matrices.
@@ 1 Introduction
The efficiency of weight vectors calculated from PCMs gained in a real-world problem was examined by Duleba and Moslem, (2019).
@@
Brunelli and Fedrizzi, (2024) formulated a property based on Pareto dominance related to the axiomatic systems used for the validation of inconsistency indices.
Szádoczki and Bozóki, (2024) examined the efficiency of weight vectors from a geometric viewpoint, and showed (mainly through examples) that the set of efficient weight vectors for a PCM with $3$ alternatives can be represented as a triangle (or a single point in the case of a consistent matrix), while for $4$ alternatives, when there are no consistent triads (triplets) among the alternatives, it can be visualized as a union of three tetrahedra. These tetrahedra are determined by the weight vectors calculated from those incomplete pairwise comparison matrices that can be obtained from the original PCM by deleting some elements, and can be represented with a line spanning tree graph. However, they did not characterize the possible cases, where the geometric properties of the set of efficient weight vectors differ, nor listed the possible cases. This paper aims to fill in this research gap, and determine the possible type of geometric characteristic of the set of efficient weight vectors, as well as show the underlying reasons, in the case of $4$ alternatives.
The rest of the paper is organized as follows. Section 2 contains the preliminaries related to pairwise comparison matrices and efficient weight vectors used throughout the study. Section 3 presents the main results, the geometric characterization of the possible cases of the set of efficient weight vectors. Finally, Section 4 concludes and raises further research questions.
@@ 2 Preliminaries
From the above definition it follows that renormalization does not affect efficiency or inefficiency.
@@
For any pairs of vertices of the BCC digraph, there is an arc in at least one of the two possible orientations, and there are arcs in both directions only if the comparison of the appropriate alternatives is estimated perfectly by the given weight vector. This latter case can only happen at the boundary of the set of efficient weight vectors (Szádoczki and Bozóki,, 2024), which is a closed set (Furtado and Johnson, 2024e, , Theorem 6), thus, the focus of this paper is the case when there is exactly one arc for each pair of vertices, i.e., the relevant BCC digraphs are tournaments.
@@
As in Furtado and Johnson, (2025), the following notations will be used in the paper. Let us denote
  • the set of efficient weight vectors for the $n\times n$ PCM $\mathbf{A}$ by $\mathcal{E}\left(\mathbf{A}\right)$;
  • the vector obtained from vector $\mathbf{w}$ by deleting (retaining) the entry indexed by $i$ as $\mathbf{w}(i)$;
  • the submatrix of matrix $\mathbf{A}$ obtained by deleting (retaining) the column and row indexed by $i$ as $\mathbf{A}(i)$;
  • the set of $\mathbf{w}$ weight vectors for which $\mathbf{w}\in\mathcal{E}\left(\mathbf{A}\right)$, and $\mathbf{w}(i)\in\mathcal{E}\left(\mathbf{A}(i)\right)$ by $\mathcal{E}\left(\mathbf{A};i\right)$.
Theorem 3 ((Furtado and Johnson,, 2025, Corollaries 13 and 14)).
Let $\mathbf{A}$ be an $n\times n$ PCM, where $n\geq 4$. Then the following hold.
  1. 1.
    $\ \mathcal{E}\left(\mathbf{A}\right)=\bigcup_{i=1}^{n}\mathcal{E}\left(\mathbf{A};i\right)=\bigcup_{i=1,\ i\neq p}^{n}\mathcal{E}\left(\mathbf{A};i\right),$
    for any $p\in\{1,\ldots,n\}$.
  2. 2.
    For any $\mathbf{w}\in\mathcal{E}\left(\mathbf{A}\right)$ there are at least two submatrices of $\mathbf{A}$, $\mathbf{A}(i)$ and $\mathbf{A}(j)$ ($i\neq j$), such that $\mathbf{w}(i)\in\mathcal{E}\left(\mathbf{A}(i)\right)$ and $\mathbf{w}(j)\in\mathcal{E}\left(\mathbf{A}(j)\right).$
This means that any $\mathbf{w}$ weight vector that is efficient for PCM $\mathbf{A}$ must be efficient for at least two submatrices of $\mathbf{A}$, where exactly one row, the corresponding column, and the related entry of the weight vector are deleted. Furthermore, the set of efficient weight vectors ($\mathcal{E}\left(\mathbf{A}\right)$) can be determined by the union of any $n-1$ of the $n$ different sets of weight vectors that are efficient for $\mathbf{A}$ and for a given submatrix of $\mathbf{A}$ ($\mathbf{A}(i)$, $i=1,\ldots,n$).
Corollary 2.
Let $\mathbf{A}$ be an $n\times n$ PCM, where $n\geq 4$, and let $\mathbf{w}$ be a weight vector for which $\mathbf{w}\in\mathcal{E}\left(\mathbf{A};i\right)$. Then all the arcs of the BCC graph $G=(V,\overrightarrow{E})_{\mathbf{A},\mathbf{w}}$ that do not include the entry indexed $i$—denoted by $G(i)=\left(V(i),\overrightarrow{E}\right)_{\mathbf{A}(i),\mathbf{w}(i)}$—are determined by $\mathbf{w}(i)$ and $\mathbf{A}(i)$, and $G(i)$ must be strongly connected, thus must have a Hamiltonian cycle.
@@ 3 Results
The goal of this paper is to determine the possible type of geometric characteristic of the set of efficient weight vectors, and to show the reasons behind that for $4$ alternatives. However, the special cases, when there are consistent triads in the PCM, are not considered, for some geometric interpretation of those, see Szádoczki and Bozóki, (2024).
@@
There is not a single consistent triad in the examined pairwise comparison matrix.
@@
$\begin{pmatrix}1&a&b&c\\ 1/a&1&d&e\\ 1/b&1/d&1&f\\ 1/c&1/e&1/f&1\end{pmatrix},$
where $a,b,c,d,e,f>0$. From now on, we use these notations.
@@
$h_{34}:\ w_{3}/w_{4}=f.$
Figure 1 shows $S_{4}$ with an example plane of $h_{12}:$ $w_{1}/w_{2}=a=2$.
@@
Figure 1: $S_{4}$ with an example plane of $h_{12}:$ $w_{1}/w_{2}=a=2$ and the related arcs of the BCC graph above and below the plane.
@@
For every element of the PCM (aside from the ones coming from the reciprocity property) there is such a plane that determines one of the arcs of the BCC graph. According to Theorem 1, if the gained BCC graph is strongly connected, then the related weight vector is efficient. The question is how can we characterize the set of efficient weight vectors in a geometric way. How many different types of sets are possible from a geometric view point, and what is the determining factor behind that?
@@
If all of the possible vectors and planes are multiplied in a given dimension by a constant, then geometrically that is only a stretching, the relation of any two vectors are not modified. This observation is formalized for the examined problem in Proposition 1.
Proposition 1.
It is sufficient to consider matrices in the form of
$\begin{pmatrix}1&1&b^{\prime}&c^{\prime}\\ 1&1&1&e^{\prime}\\ 1/b^{\prime}&1&1&1\\ 1/c^{\prime}&1/e^{\prime}&1&1\end{pmatrix}.$
Proof.
Let us consider
$\mathbf{A}=\begin{pmatrix}1&a&b&c\\ 1/a&1&d&e\\ 1/b&1/d&1&f\\ 1/c&1/e&1/f&1\end{pmatrix},\mathbf{w}=\begin{pmatrix}w_{1}\\ w_{2}\\ w_{3}\\ w_{4}\end{pmatrix},$
where $\mathbf{w}$ is the weight vector obtained from $\mathbf{A}$ with any chosen weight vector calculation method. Let us multiply the weight of the first alternative with $1/adf$ (a shorter notation for $1/(adf)$). This is akin to a measurement unit change. This will result in a corresponding change in $\mathbf{A}$. We get
$\begin{pmatrix}1&1/df&b/adf&c/adf\\ df&1&d&e\\ adf/b&1/d&1&f\\ adf/c&1/e&1/f&1\end{pmatrix},\begin{pmatrix}w_{1}/adf\\ w_{2}\\ w_{3}\\ w_{4}\end{pmatrix}.$
Now let us multiply the second alternative with $1/df$. We get
$\begin{pmatrix}1&1&b/adf&c/adf\\ 1&1&1/f&e/df\\ adf/b&f&1&f\\ adf/c&df/e&1/f&1\end{pmatrix},\begin{pmatrix}w_{1}/adf\\ w_{2}/df\\ w_{3}\\ w_{4}\end{pmatrix}.$
Now let us multiply the third alternative with $1/f$. We get
$\mathbf{A}^{\prime}=\begin{pmatrix}1&1&b/ad&c/adf\\ 1&1&1&e/df\\ ad/b&1&1&1\\ adf/c&df/e&1&1\end{pmatrix},\mathbf{w}^{\prime}=\begin{pmatrix}w_{1}/adf\\ w_{2}/df\\ w_{3}/f\\ w_{4}\end{pmatrix}.$
Thus, with $b^{\prime}=b/ad,c^{\prime}=c/adf,e^{\prime}=e/df$ the matrix has the required form. ∎
Note that after the transformation $\sum w_{i}=1$ no longer holds. This means that instead of $S_{4}$, the weight vectors should be illustrated in a tetrahedron, where the vertices are the following:
$(w_{1}=1/adf,w_{2}=0,w_{3}=0,w_{4}=0),$
$(w_{1}=0,w_{2}=1/df,w_{3}=0,w_{4}=0),$
$(w_{1}=0,w_{2}=0,w_{3}=1/f,w_{4}=0),$
$(w_{1}=0,w_{2}=0,w_{3}=0,w_{4}=1).$
However, the geometric nature of the set of efficient weight vectors will be the same as the one for the matrices that follow the form presented in Proposition 1 in $S_{4}$. Thus, we will only focus on those to determine the different types of geometric characteristic of sets of efficient weight vectors.
It is also true that we can get the same geometric object with reversed dimensions (looking at it from different angles). This is formalized for pairwise comparison matrices in Proposition 2.
@@
$\mathbf{A}=\begin{pmatrix}1&a&b&c\\ 1/a&1&d&e\\ 1/b&1/d&1&f\\ 1/c&1/e&1/f&1\end{pmatrix}$
can be transformed with appropriate reindexing of alternatives to the following form:
$\mathbf{A}^{\prime\prime}=\begin{pmatrix}1&a^{\prime\prime}&b^{\prime\prime}&c^{\prime\prime}\\ 1/a^{\prime\prime}&1&d^{\prime\prime}&e^{\prime\prime}\\ 1/b^{\prime\prime}&1/d^{\prime\prime}&1&f^{\prime\prime}\\ 1/c^{\prime\prime}&1/e^{\prime\prime}&1/f^{\prime\prime}&1\end{pmatrix},$
where $\mathbf{A}^{\prime\prime}$ falls into one of the following cases:
Case 1: $a^{\prime\prime}d^{\prime\prime}<b^{\prime\prime}$, $d^{\prime\prime}f^{\prime\prime}<e^{\prime\prime}$, $b^{\prime\prime}f^{\prime\prime}<c^{\prime\prime}$, $a^{\prime\prime}e^{\prime\prime}<c^{\prime\prime}$.
Case 2: $a^{\prime\prime}d^{\prime\prime}<b^{\prime\prime}$, $d^{\prime\prime}f^{\prime\prime}<e^{\prime\prime}$, $b^{\prime\prime}f^{\prime\prime}<c^{\prime\prime}$, $a^{\prime\prime}e^{\prime\prime}>c^{\prime\prime}$.

### v2

@@ Abstract
Pairwise comparison matrices and the weight vectors obtained from them are important concepts in multi-criteria decision making. A weight vector calculated from a pairwise comparison matrix is called Pareto efficient if the approximation of the matrix elements by the weight ratios cannot be improved for any element of the matrix without worsening it for another element. The aim of this paper is to show the geometrical properties of the set of Pareto efficient weight vectors for $4\times 4$ pairwise comparison matrices. We prove that the set of efficient weight vectors is a union of three tetrahedra, each determined by four weight vectors calculated from an incomplete submatrix of the pairwise comparison matrix that can be represented by a path-type spanning tree graph. It is shown that with suitable rearrangements the orientations of the $4$-cycles in the Blanquero-Carrizosa-Conde graphs for the efficient weight vectors are determined as well. The special cases (double perturbed, simple perturbed, consistent matrices) are discussed in the appendices.
@@ 1 Introduction
The efficiency of weight vectors calculated from PCMs related to a real-world problem was examined by Duleba and Moslem, (2019).
@@
Brunelli and Fedrizzi, (2024) formulated a property related to the axiomatic systems used for the validation of inconsistency indices based on Pareto dominance.
Szádoczki and Bozóki, (2024) examined the efficiency of weight vectors from a geometric viewpoint, and showed (mainly through examples) that the set of efficient weight vectors for a PCM with $3$ alternatives can be represented as a triangle (or a single point in the case of a consistent matrix), while for $4$ alternatives, when there are no consistent triads (triplets) among the alternatives, it can be visualized as a union of three tetrahedra. These tetrahedra are determined by the weight vectors calculated from those incomplete pairwise comparison matrices that can be obtained from the original PCM by deleting some elements, and can be represented with a path-type spanning tree graph. However, they did not prove the generality of their findings, did not characterize the possible cases, where the geometric properties of the set of efficient weight vectors differ, nor listed the possible cases, only presented different examples. This paper aims to fill in this research gap, and determine the possible type of geometric characteristic of the set of efficient weight vectors, as well as show the underlying reasons, in the case of $4$ alternatives.
The rest of the paper is organized as follows. Section 2 contains the preliminaries related to pairwise comparison matrices and efficient weight vectors used throughout the study. Section 3 presents the main results, the geometric characterization of the possible cases of the set of efficient weight vectors. Finally, Section 4 concludes and raises further research questions. The results of Section 3 are given with more examples and including special cases as well (double perturbed, simple perturbed, consistent matrices) in Appendices Appendix A and Appendix B.
@@ 2 Preliminaries
It may happen that there are missing elements in the PCM. That is called an incomplete pairwise comparison matrix (IPCM).
Definition 3 (Incomplete pairwise comparison matrix (IPCM)).
An $n\times n$ matrix $\mathbf{A}=[a_{ij}]$ is an incomplete pairwise comparison matrix (IPCM) if:
[…]
    • $a_{ji}=1/a_{ij}$ if $a_{ij}\in\mathbb{R}_{+}$,
    • $a_{ji}=\ast$ if $a_{ij}=\ast$,
where $\ast$ denotes the missing elements, and $\mathbb{R}_{+}$ is the set of positive real numbers.
Most of the properties of IPCMs can be suitably handled by their representing graphs (graphs of comparisons).
Definition 4 (Representing graph/Graph of comparisons).
An incomplete pairwise comparison matrix $\mathbf{A}$ can be represented by an undirected graph $G=(V,E)$, where:
  • the vertices $V=\{1,2,\ldots,n\}$ correspond to the alternatives,
  • while the edge set $E$ represents the known elements of $\mathbf{A}$ outside the main diagonal:
    $e_{ij}\in E\iff a_{ij}\neq\ast\ \text{and}\ i\neq j.$
One can calculate a weight vector from an IPCM as well, where the different methods rely on complementing the matrix as consistently as possible (Tekile et al.,, 2023). In order to calculate a unique weight vector from an $n\times n$ IPCM, it is necessary that the matrix has a connected representing graph (Bozóki et al.,, 2010). To reach this, at least $n-1$ comparisons (edges) are needed. Thus, the smallest units from which a unique weight vector can be calculated are the IPCMs that can be represented by a spanning tree (see Example 1), which can always be complemented perfectly consistently. These units also have an important role in determining the set of Pareto efficient weight vectors.
Example 1 (Spanning tree representing graph).
Let us consider the following incomplete pairwise comparison matrix:
$\mathbf{A}=\begin{pmatrix}1&a_{12}&\ast&\ast\\ 1/a_{12}&1&a_{23}&\ast\\ \ast&1/a_{23}&1&a_{34}\\ \ast&\ast&1/a_{34}&1\end{pmatrix}$
The representing graph $G$ of $\mathbf{A}$ is shown in Figure 1.
[…]
Figure 1: The representing graph of $\mathbf{A}$.
The exact values of the known comparisons do not matter regarding the representing graph of an IPCM.
@@
From the above definition it follows that renormalization does not affect efficiency or inefficiency, thus we focus on the case, when the weight vector is normalized ($\sum_{i=1}^{n}w_{i}=1$).
@@
Example 2 (BCC directed graph).
Let us consider the following pairwise comparison matrix $\mathbf{A}$, and weight vector $\mathbf{w}$:
$\mathbf{A}=\begin{pmatrix}1&a_{12}&a_{13}&a_{14}\\ 1/a_{12}&1&a_{23}&a_{24}\\ 1/a_{13}&1/a_{23}&1&a_{34}\\ 1/a_{14}&1/a_{24}&1/a_{34}&1\end{pmatrix}=\begin{pmatrix}1&1&5&7\\ 1&1&2&8\\ 1/5&1/2&1&1/3\\ 1/7&1/8&3&1\end{pmatrix},$
$w=\left(0.25,0.25,0.25,0.25\right)^{T}.$
The related $G=(V,\overrightarrow{E})_{\mathbf{A},\mathbf{w}}$ BCC directed graph can be seen in Figure 2.
1234
Figure 2: The BCC digraph related to $\mathbf{A}$ and $\mathbf{w}$.
It can be seen that for the comparison between alternatives $1$ and $2$, the estimation is perfect ($w_{1}/w_{2}=a_{12}=1$), thus there are arcs between them with both orientations. However, $G$ is not strongly connected, thus according to Theorem 1, $\mathbf{w}$ is not an efficient weight vector of $\mathbf{A}$.
For any pairs of vertices of the BCC digraph, there is an arc in at least one of the two possible orientations, and there are arcs in both directions only if the comparison of the appropriate alternatives is estimated perfectly by the given weight vector. To determine the set of efficient weight vectors, it is enough to focus on the case when there is exactly one arc for each pair of vertices, i.e., the relevant BCC digraphs are tournaments.
@@
In the case when some of the elements are estimated perfectly, there still must be at least one Hamiltonian cycle in the associated BCC diagraph, as every arc has at least one orientation.
@@ 3 Results
The goal of this paper is to determine the possible type of geometric characteristic of the set of efficient weight vectors, and to show the reasons behind that for $4$ alternatives. However, the special cases, when there are consistent cycles in the PCM, are only considered in Appendix Appendix B, for some geometric interpretation of some of those instances, see also Szádoczki and Bozóki, (2024).
@@
The examined pairwise comparison matrix is a triple perturbed PCM, it has no consistent $k$-cycles ($k\geq 3$).
@@
$\begin{pmatrix}1&a_{12}&a_{13}&a_{14}\\ 1/a_{12}&1&a_{23}&a_{24}\\ 1/a_{13}&1/a_{23}&1&a_{34}\\ 1/a_{14}&1/a_{24}&1/a_{34}&1\end{pmatrix},$
where $a_{12},a_{13},a_{14},a_{23},a_{24},a_{34}>0$. From now on, we use these notations.
@@
$h_{34}:\ w_{3}/w_{4}=a_{34}.$
Figure 3 shows $S_{4}$ with an example plane of $h_{12}:$ $w_{1}/w_{2}=a_{12}=2$.
@@
Figure 3: $S_{4}$ with an example plane of $h_{12}:$ $w_{1}/w_{2}=a_{12}=2$ and the related arcs of the BCC graph above and below the plane.
@@
For every element of the PCM (aside from the ones coming from the reciprocity property) there is such a plane that determines one of the arcs of the BCC graph. According to Theorem 1, if the gained BCC graph is strongly connected, then the related weight vector is efficient.
Szádoczki and Bozóki, (2024) conjectured via an example that the set of efficient weight vectors for $4$ alternatives geometrically consists of the union of three tetrahedra (denoted by $\mathcal{T}_{1},\mathcal{T}_{2},$ and $\mathcal{T}_{3}$) associated with the labeled Hamiltonian cycles of BCC graphs presented in Figure 4, where for a given PCM always only one of the two directions is possible for each cycle. This conjecture is proved in the present work.
[…]
Figure 4: The labeled Hamiltonian cycles of BCC graphs associated with tetrahedra $\mathcal{T}_{1},\mathcal{T}_{2},$ and $\mathcal{T}_{3}$. Each tetrahedron is related to the cycles below them, from which always only one is possible for a given PCM (see Lemma 2).
Each of $\mathcal{T}_{1},\mathcal{T}_{2},$ and $\mathcal{T}_{3}$ is determined by exactly four of the six $h_{ij}$ planes ($i=1,\ldots,3,\ j=i+1,\ldots,4$).
The question is, which combinations of the Hamiltonian cycles associated with $\mathcal{T}_{1},\mathcal{T}_{2},$ and $\mathcal{T}_{3}$ are possible. What are the determining factors behind that, and how can we describe that in a geometric manner?
@@
It is true that we can get the same geometric object with reversed dimensions (looking at it from different angles), and the properties of pairwise comparisons are often examined through their triads (the different triplets of alternatives) as their building blocks. However, in the examined problem it is not enough to focus on the triads.
Remark 3.
The properties of the set of efficient weight vectors is not solely dependent on the triads of a $4\times 4$ pairwise comparison matrix. Furthermore, such a matrix can be a double perturbed one (with consistent $4$-cycles violating Assumption 1) even in the case when all of its triads are inconsistent. For further details, see Appendix Appendix A.
Instead of the triads, the 4-cycles should be considered, for which the dimension reversion is formalized in Proposition 1.
@@
$\mathbf{A}=\begin{pmatrix}1&a_{12}&a_{13}&a_{14}\\ 1/a_{12}&1&a_{23}&a_{24}\\ 1/a_{13}&1/a_{23}&1&a_{34}\\ 1/a_{14}&1/a_{24}&1/a_{34}&1\end{pmatrix},$
for which Assumption 1 holds, can be transformed with appropriate reindexing of alternatives to the following form:
$\mathbf{A}^{\prime\prime}=\begin{pmatrix}1&a_{12}^{\prime\prime}&a_{13}^{\prime\prime}&a_{14}^{\prime\prime}\\ 1/a_{12}^{\prime\prime}&1&a_{23}^{\prime\prime}&a_{24}^{\prime\prime}\\ 1/a_{13}^{\prime\prime}&1/a_{23}^{\prime\prime}&1&a_{34}^{\prime\prime}\\ 1/a_{14}^{\prime\prime}&1/a_{24}^{\prime\prime}&1/a_{34}^{\prime\prime}&1\end{pmatrix},$
where $\mathbf{A}^{\prime\prime}$ has the following property:
$a_{12}^{\prime\prime}a_{23}^{\prime\prime}a_{34}^{\prime\prime}a_{41}^{\prime\prime}<1$,  $a_{14}^{\prime\prime}a_{42}^{\prime\prime}a_{23}^{\prime\prime}a_{31}^{\prime\prime}<1$,  $a_{13}^{\prime\prime}a_{34}^{\prime\prime}a_{42}^{\prime\prime}a_{21}^{\prime\prime}<1$.
where the notation $a_{ij}$ with $i>j$ stands for $1/a_{ji}$.